A tennis ball can in the shape of a cylinder with a flat top and bottom of the same radius as the tennis balls is designed so the space inside the can that is not occupied by the balls has a volume at most equal to the volume of one ball. What is the largest number of balls that the can will contain?
Answer by edjones(8007)

Cylinder A=pi*r^2*h
Sphere A= 4/3*pi*r^3
.
for 1 ball with a radius of 1
Cylinder A=6/3*pi (h=2r=6/3)
Sphere A= 4/3*pi
------------------
difference 2/3*pi
.
for 2 ball with a radius of 1
Cylinder A=12/3*pi
Sphere A= 8/3*pi
------------------
difference 4/3*pi= volume of 1 ball
.
Can will hold no more than 2 balls.
Ed